# Request a Note: How the Request Function Shapes X’s Community Notes System

### v1

@@ Abstract.
X’s Community Notes is a crowdsourced fact-checking system. To improve its scalability, X recently introduced “Request Community Note” feature, enabling users to solicit fact-checks from contributors on specific posts. Yet, its implications for the system—what gets checked, by whom, and with what quality—remain unclear. Using $98,685$ requested posts and their associated notes, we evaluate how requests shape the Community Notes system. We find that contributors prioritize posts with higher misleadingness and from authors with greater misinformation exposure, but neglect political content emphasized by requestors. Selection also diverges along partisan lines: contributors more often annotate posts from Republicans, while requestors surface more from Democrats. Although only 12% of posts receive request-fostered notes from top contributors, these notes are rated as more helpful and less polarized than others, partly reflecting top contributors’ selective fact-checking of misleading posts. Our findings highlight both the limitations and promise of requests for scaling high-quality community-based fact-checking.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Social mediaccs: Information systems Crowdsourcingccs: Human-centered computing Empirical studies in collaborative and social computingccs: Security and privacy Human and societal aspects of security and privacy
[…]
@@ 1. Introduction
Designing effective countermeasures to identify falsehoods and curb the spread of misleading posts is still a challenge. Years ago, the social media platform $\mathbb{X}$ (formerly Twitter) adopted Community Notes, a crowdsourced fact-checking system designed to address misleading posts and provide users with helpful additional context. This initiative marks the first large-scale use of community-based fact-checking by a major social media platform (Chuai et al., 2024b; Augenstein et al., 2025). A subset of users—called contributors—are periodically admitted into the program if they meet specific eligibility requirements, such as having no recent violations of $\mathbb{X}$’s rules. Contributors begin by rating the helpfulness of existing notes, and only after demonstrating reliability can they unlock the ability to write notes themselves. All notes are evaluated by a note selection algorithm, and then those deemed helpful by the algorithm are displayed directly on the relevant posts (Notes, 2025f). Inspired by this program, other social media providers—including YouTube, Meta, and TikTok—are beginning to explore similar community-based moderation models (Presser, 2025; YouTube, 2024; Kaplan, 2025). For example, Meta is piloting the Community Notes program across Facebook, Instagram, and Threads in the U.S., while TikTok plans to offer a similar feature called Footnotes to provide context and corrections for misleading posts. Given the rapid expansion of community-based fact-checking system, it is crucial to understand how the system functions in the real world.
Recent work has shown that $\mathbb{X}$’s Community Notes program can reliably attach helpful notes to the corresponding misleading posts (Wojcik et al., 2022; Allen et al., 2024). Moreover, community fact-checks are perceived as more trustworthy than expert fact-checks, can effectively reduce the spread of misleading posts, and even prompt authors to delete their problematic content (Drolsbach et al., 2024; Chuai et al., 2024a; Wojcik et al., 2022; Chuai et al., 2025a). However, the Community Notes system, in its current implementation on $\mathbb{X}$, still faces limitations, such as the small number of displayed notes relative to all generated notes and restrictions on who can author them (Chuai et al., 2024a).
@@
Research questions. In this study, we examine the newly introduced Request Community Note feature on $\mathbb{X}$, tracing the process from the submission of requests to the evaluation of the associated notes. The workflow involves requestors flagging posts (i. e., requested posts) for note writers, note writers selecting requested posts to propose community notes, and the note selection algorithm evaluating the generated notes. Notably, note writers retain autonomy: they can independently choose any post to fact-check and propose community notes, regardless of whether they see request alerts or not. Moreover, the note selection algorithm evaluates community notes along two key dimensions: helpfulness, indicating the consensus of perceived helpfulness among heterogeneous contributors, and polarization, capturing the degree of variation in ratings. Accordingly, we organize our investigation around three research questions to explore how note writers select requested posts to fact-check, the extent to which their selections are influenced by requests, and how community notes fostered by user requests (i. e., request-fostered notes) differ from other notes in helpfulness and polarization according to the note selection algorithm:
@@
Contributions. Overall, our study results in three main findings. (i) The likelihood that a requested post receives a community note is associated with the post’s content and its author. For instance, posts related to entertainment or finance, with higher estimated misleadingness, or from accounts with higher misinformation exposure scores, are more likely to receive community notes. Our findings suggest that Community Notes contributors and requestors may have different selection patterns. (ii) With more than half (53.6%) of requested posts receiving community notes, only an estimated 12.1% of those are likely fostered by user requests, showing that the effect of requesting community notes in fostering note generation is, so far, limited. Nonetheless, requests tend to shift contributors’ attention toward political content, while still prioritizing posts with higher misleadingness (as measured by GPT-4.1 from OpenAI). (iii) Request-fostered notes generated by top writers are evaluated as more helpful and less polarized than other notes, a pattern that likely reflects writers’ selective prioritization of requested posts with higher levels of misleadingness. To our knowledge, our study provides the first empirical evaluation of this novel request mechanism, offering insights into both the opportunities and challenges of scaling crowdsourced fact-checking systems.
@@ 2.1. Misinformation Identification and Fact-Checking
To identify misinformation, traditional human-driven fact-checking often replies on professional experts and third-party organizations. However, the scarcity of expert fact-checkers, coupled with the relentless pace of online content, hampers the timely and scalable implementation of professional fact-checking (Chuai et al., 2024b; Augenstein et al., 2025). Furthermore, despite its depth and rigor, professional fact-checking has faced criticism, as experts are sometimes perceived as biased in selecting which claims to review—raising concerns about agenda-setting and potentially eroding public trust in their assessments (Chuai et al., 2025b; Lee et al., 2023). To address the limitations of professional fact-checking, crowdsourced (or community-based) fact-checking approaches have been proposed, leveraging the wisdom of crowds to achieve broader and faster coverage of online content (Allen et al., 2021). Although individual assessments may be subject to bias and noise (Allen et al., 2022), aggregated judgments—even from relatively small crowds—have been shown to be reliable and comparable in accuracy to expert evaluations (Bhuiyan et al., 2020; Epstein et al., 2020; Pennycook and Rand, 2019; Saeed et al., 2022). Moreover, crowdsourced assessments have the potential to mitigate public trust issues associated with expert fact-checks (Allen et al., 2021; Zhang et al., 2024; Drolsbach et al., 2024).
@@
Recent advances in Large Language Models (LLMs), such as OpenAI’s GPT series, introduce novel opportunities to overcome the shortcomings of traditional machine learning models in fact-checking. LLMs exhibit strong contextual understanding and can simulate human-like reasoning processes, allowing them to evaluate claims not only on textual cues but also in relation to broader world knowledge. They can flexibly incorporate external evidence through retrieval-augmented generation, adapt to diverse domains with minimal task-specific training, and generate explanations that improve transparency in the fact-checking process (Chen and Shu, 2024; Warren et al., 2025). Previous work suggests that LLMs can achieve human-comparable performance in misinformation detection tasks (Choi and Ferrara, 2024; DeVerna et al., 2024; Costabile et al., 2025). For instance, ChatGPT has been shown to achieve an accuracy of approximately 90% in identifying false headlines, demonstrating the potential of LLMs to support large-scale misinformation detection (DeVerna et al., 2024). At the same time, LLM-based fact-checking is not without challenges. LLMs tend to adopt a more structured evaluation strategy, whereas human annotators often display greater variability in their use of evaluative criteria, particularly for borderline or ambiguous claims (Costabile et al., 2025). More critically, their influence on user perception can be harmful: when LLMs mislabel true claims as false, they risk reducing belief in accurate information; conversely, when they express uncertainty about false claims, they may inadvertently increase belief in misinformation (DeVerna et al., 2024; Lu et al., 2025). These limitations underscore that, while LLMs perform strongly in detecting misinformation, human judgment remains essential for deciding what content to surface to users, helping to mitigate unintended harms.
@@ 2.2. Crowdsourced Fact-Checking in Practice: X’s Community Notes
Given the necessity of human-centered approaches and the demonstrated promise of crowdsourced fact-checking, $\mathbb{X}$ introduced Community Notes, a system that engages diverse users in collaboratively annotating potentially misleading content at scale (Notes, 2025f). First, Community Notes contributors can append contextual information or corrections to potentially misleading posts, offering alternative perspectives or clarifications. Contributors are required to cite external sources in their community notes. These notes are continuously evaluated through peer ratings from other contributors. Additionally, $\mathbb{X}$ develops a bridging algorithm to feature objectively informative community notes that are subjectively perceived to be helpful across heterogeneous contributors (Wojcik et al., 2022). Beyond contributors, all users can now participate indirectly through the recently launched Request Community Note function, which allows any account with a verified phone number to submit requests for posts they believe would benefit from a community note. When enough requests are submitted, top writers can see alerts highlighting these posts and can choose to write a note in response (see the illustrations in Fig. 1). This division of roles—between note writers, note raters, and requestors—illustrates how Community Notes distributes responsibilities across different layers of the user base, thereby expanding participation beyond the core contributor group.
Research has shown that Community Notes system can successfully identify misleading content, reduce its spread, and even pressure authors to delete problematic posts (Wojcik et al., 2022; Martel and Rand, 2024; Allen et al., 2024; Chuai et al., 2024a). Additionally, community-based annotations are often perceived as more trustworthy than professional fact-checks (Drolsbach et al., 2024). These encouraging findings suggest that Community Notes is promising in addressing misinformation on social media platforms and have sparked growing scholarly interest in understanding the dynamics of the system and identifying opportunities to refine and enhance its design (Augenstein et al., 2025). For instance, despite the relatively large volume of notes generated by contributors, only a small fraction—around 10%—achieve sufficient support from diverse users and are displayed to users (De et al., 2025; Chuai et al., 2024a). At the same time, the display process is often not fast enough to intervene before misleading posts have already begun to spread widely (Chuai et al., 2024a, b). To address these limitations, recent work has explored two complementary directions.
@@ 3.1. Data Collection
Subsequently, we used the full-archive search endpoint of the $\mathbb{X}$ Pro API to collect the requested posts. In total, the submitted requests target $2,427,451$ unique posts. However, according to the platform guidelines at the time of data collection, a post must receive at least 5 requests to be surfaced to top writers, indicating a threshold for which posts are eligible for community evaluation (Notes, 2025e). In our dataset, $154,090$ posts meet this criterion, representing only 6.3% of all requested posts. However, the requests submitted to these posts account for more than half (52.6%) of the total requests. This suggests that, while users submitted a large volume of requests, attention was concentrated on a small set of posts. Based on the corresponding IDs of these eligible posts, we successfully retrieved $133,351$ posts, of which $98,685$ (74%) are in English and were authored by $22,915$ unique accounts. We restrict our analysis to the collected posts in English and focus on the request feature in the U.S. context. Next, we introduce how to extract the characteristics of posts and their authors for our subsequent analysis (see summary statistics in Table 1).
Table 1. Overview of requested posts. The three columns are for all requested posts, posts with community notes, and posts without community notes, respectively. Reported are mean values or count numbers for the variables (standard deviations in parentheses).
@@
   Content type: Claim 0.666 (0.309) 0.664 (0.315) 0.667 (0.302)
   Content type: Opinion 0.522 (0.335) 0.507 (0.336) 0.540 (0.333)
   GPT misleadingness 0.428 (0.317) 0.435 (0.323) 0.421 (0.310)
@@
   Account type: Blue 72.5% 72.3% 72.6%
   Account type: Business 6.6% 6.4% 6.7%
   Account type: Government 7.2% 6.5% 7.9%

### v2

# arXiv:2509.09956v2[cs.SI] 02 Mar 2026

@@ Abstract.
$\mathbb{X}$’s Community Notes is a crowdsourced fact-checking system. To improve its scalability, $\mathbb{X}$ introduced “Request Community Note” feature, enabling users to solicit fact-checks from contributors on specific posts. Yet, its implications for the system—what gets checked, by whom, and with what quality—remain unclear. Using $98,685$ requested posts and their associated notes, we evaluate how requests shape the Community Notes system. We find that requested posts with higher GPT-estimated misleadingness and from authors with greater misinformation exposure are more likely to receive notes. Conversely, requested political posts (vs. non-political) are less likely to receive notes. We also observe partisan asymmetries: posts from Republicans are more likely to receive notes than those from Democrats. Although only 12% of requested posts receive request-fostered notes from top contributors, these notes are rated as more helpful and less polarized than others, partly reflecting top contributors’ selective fact-checking of misleading posts. Our findings highlight both the limitations and promise of requests for scaling high-quality community-based fact-checking.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790524isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Social mediaccs: Information systems Crowdsourcingccs: Human-centered computing Empirical studies in collaborative and social computingccs: Security and privacy Human and societal aspects of security and privacy
@@
The introduction of ”Request Community Note” feature on X, sourced from the official Community Notes website. It allows users to request a note on a post they believe would benefit from additional context (see the left part of the figure). When a sufficient number of requests are submitted, a subset of Community Notes contributors (i.e., top writers) will see an alert below the corresponding requested post, and can choose to propose a note (see the right part of the figure).
@@ 1. Introduction
Designing effective countermeasures to identify falsehoods and curb the spread of misleading posts (i. e., posts carrying misinformation) on social media platforms is still a challenge. Years ago, the social media platform $\mathbb{X}$ (formerly Twitter) adopted Community Notes, a crowdsourced fact-checking system designed to address misleading posts and provide users with helpful additional context. [^1] Here, misinformation refers to false or misleading information that is contradicted by empirical evidence or inconsistent with common or expert consensus, without imposing assumptions on the intent or the format of the content (Chuai et al., 2024b). It serves as an umbrella term including not only factually incorrect content but also disinformation and fake news—both deliberately fabricated to mislead others and/or cause harms (Pennycook and Rand, 2021; Chuai et al., 2024b; Ecker et al., 2022). Prior work further shows that factually accurate yet misleading content can be more influential than outright falsehoods (Allen et al., 2024a). In the Community Notes system, the criteria used for labeling misleading posts align closely with our definition of misinformation (see details in Suppl. S1). This initiative marks the first large-scale implementation of community-based fact-checking by a major social media platform (Chuai et al., 2024b; Augenstein et al., 2025). A subset of users—called contributors—are periodically admitted into the program if they meet specific eligibility requirements, such as having no recent violations of $\mathbb{X}$’s rules. Contributors begin by rating the helpfulness of existing notes, and only after demonstrating reliability can they unlock the ability to write notes themselves. All notes are evaluated by a note selection algorithm, and then those deemed helpful by the algorithm are displayed directly on the relevant posts (Notes, 2025f). Inspired by this program, other social media providers, including YouTube, Meta, and TikTok, are beginning to explore similar community-based moderation models (Presser, 2025; YouTube, 2024; Kaplan, 2025). For example, Meta is piloting the Community Notes program across Facebook, Instagram, and Threads in the U.S., while TikTok plans to offer a similar feature called Footnotes to provide context and corrections for misleading posts. Given the rapid expansion of community-based fact-checking system, it is crucial to understand how the system functions in the real world.
Recent work has shown that $\mathbb{X}$’s Community Notes program can reliably attach helpful notes to the corresponding misleading posts (Wojcik et al., 2022; Allen et al., 2024b). Moreover, community fact-checks are perceived as more trustworthy than expert fact-checks, can effectively reduce the spread of misleading posts, and even prompt authors to delete their problematic content (Drolsbach et al., 2024; Chuai et al., 2024a; Wojcik et al., 2022; Chuai et al., 2025a). However, the Community Notes system, in its current implementation on $\mathbb{X}$, still faces limitations, such as the small number of displayed notes relative to all generated notes and restrictions on who can author them (Chuai et al., 2024a).
@@
Research questions. In this study, we examine the newly introduced Request Community Note feature on $\mathbb{X}$, tracing the process from the submission of requests to the evaluation of the associated notes. The workflow involves requestors flagging posts (i. e., requested posts) for note writers, note writers selecting requested posts to propose community notes, and the note selection algorithm evaluating the generated notes. Notably, note writers retain autonomy: they can independently choose any post to fact-check and propose community notes, regardless of whether they see request alerts or not. Moreover, the note selection algorithm evaluates community notes along two key dimensions: helpfulness, indicating the consensus of perceived helpfulness among heterogeneous contributors, and polarization, capturing the degree of variation in ratings. Accordingly, we organize our investigation around three research questions (RQs) to explore how note writers select requested posts to fact-check, the extent to which their selections are influenced by requests, and how community notes fostered by user requests (i. e., request-fostered notes) differ from other notes in helpfulness and polarization according to the note selection algorithm:
@@
Contributions. Overall, our study results in three main findings. (i) The likelihood that a requested post receives a community note is associated with the post’s content and its author. For instance, posts related to entertainment or finance, containing claim-oriented content, with higher estimated misleadingness, or from accounts with higher misinformation exposure scores, are more likely to receive community notes. Our findings suggest that Community Notes contributors and requestors may have different selection patterns. (ii) With more than half (53.6%) of requested posts receiving community notes, only an estimated 12.1% of those are likely fostered by user requests, showing that the effect of requesting community notes in fostering note generation is, so far, limited. Nonetheless, requests tend to shift contributors’ attention toward political content, while prioritizing posts with higher misleadingness (as measured by GPT-4.1 from OpenAI). (iii) Request-fostered notes generated by top writers are evaluated as more helpful and less polarized than other notes, a pattern that likely reflects writers’ selective prioritization of requested posts with higher levels of misleadingness. To our knowledge, our study provides the first empirical evaluation of this novel request mechanism, offering insights into both the opportunities and challenges of scaling crowdsourced fact-checking systems.
@@ 2.1. Misinformation Identification and Fact-Checking
To identify misinformation, traditional human-driven fact-checking often relies on professional experts and third-party organizations. However, the scarcity of expert fact-checkers, coupled with the relentless pace of online content, hampers the timely and scalable implementation of professional fact-checking (Chuai et al., 2024b; Augenstein et al., 2025). Furthermore, despite its depth and rigor, professional fact-checking has faced criticism, as experts are sometimes perceived as biased in selecting which claims to review—raising concerns about agenda-setting and potentially eroding public trust in their assessments (Chuai et al., 2025b; Lee et al., 2023). To address the limitations of professional fact-checking, crowdsourced (or community-based) fact-checking approaches have been proposed, leveraging the wisdom of crowds to achieve broader and faster coverage of online content (Allen et al., 2021; Godel et al., 2021). Early explorations of crowdsourced verification demonstrate that journalists and online volunteer communities can collaboratively assess information accuracy and provide real-time event coverage during emergencies (Dailey and Starbird, 2014). Furthermore, hybrid crowdsourcing processes that involve both local and remote participants help enhance the completeness of information (Agapie et al., 2015). Although individual assessments may be subject to bias and noise (Allen et al., 2022), aggregated judgments—even from relatively small crowds—have been shown to be reliable and comparable in accuracy to expert evaluations (Bhuiyan et al., 2020; Epstein et al., 2020; Pennycook and Rand, 2019; Saeed et al., 2022; Allen et al., 2021). Moreover, crowdsourced assessments have the potential to mitigate public trust issues associated with expert fact-checks (Zhang et al., 2024; Drolsbach et al., 2024; Chuai et al., 2025b). Together, these studies provide important foundations for the development of community-based fact-checking systems.
@@
Recent advances in Large Language Models (LLMs), such as OpenAI’s GPT series, introduce novel opportunities to overcome the shortcomings of traditional machine learning models in fact-checking. LLMs exhibit strong contextual understanding and can simulate human-like reasoning processes, allowing them to evaluate claims not only on textual cues but also in relation to broader world knowledge. They can flexibly incorporate external evidence through retrieval-augmented generation, adapt to diverse domains with minimal task-specific training, and generate explanations that improve transparency in the fact-checking process (Chen and Shu, 2024; Warren et al., 2025). Previous work suggests that LLMs can achieve human-comparable performance in misinformation detection tasks (Choi and Ferrara, 2024; DeVerna et al., 2024; Costabile et al., 2025; Rathje et al., 2024; Majer and Šnajder, 2024). For instance, ChatGPT has been shown to achieve an accuracy of approximately 90% in identifying false headlines, demonstrating the potential of LLMs to support large-scale misinformation detection (DeVerna et al., 2024). At the same time, LLM-based fact-checking is not without challenges. LLMs tend to adopt a more structured evaluation strategy, whereas human annotators often display greater variability in their use of evaluative criteria, particularly for borderline or ambiguous claims (Costabile et al., 2025). More critically, their influence on user perception can be harmful: when LLMs mislabel true claims as false, they risk reducing belief in accurate information; conversely, when they express uncertainty about false claims, they may inadvertently increase belief in misinformation (DeVerna et al., 2024; Lu et al., 2025). These limitations underscore that, while LLMs can perform well in detecting misinformation, human judgment remains essential for deciding what content to surface to users, helping to mitigate unintended harms.
@@ 2.2. Crowdsourced Fact-Checking in Practice: X’s Community Notes
Given the necessity of human-centered approaches and the demonstrated promise of crowdsourced fact-checking, $\mathbb{X}$ introduced Community Notes, a system that engages diverse users in collaboratively annotating potentially misleading content at scale (Notes, 2025f). Specifically, Community Notes contributors can append contextual information or corrections to potentially misleading posts, offering alternative perspectives or clarifications. Contributors are required to cite external sources in their community notes. These notes are continuously evaluated through peer ratings from other contributors. However, featuring community notes based on majority-based aggregation rules makes the system vulnerable to manipulation or one-side ratings from ideologically homogeneous raters. Research shows that politically balanced groups produce more reliable and accurate evaluations than homogeneous ones (Allen et al., 2021; Godel et al., 2021). Therefore, $\mathbb{X}$ develops a bridging algorithm to feature objectively informative community notes that are subjectively perceived to be helpful across heterogeneous contributors (Wojcik et al., 2022). Beyond contributors, prior research has also explored on-demand fact-checking models that enable ordinary users to request fact-checks for content they wish to have checked (Kriplean et al., 2014). In line with this idea, all users on $\mathbb{X}$ can now participate indirectly through the recently launched Request Community Note function, which allows any account with a verified phone number to submit requests for posts they believe would benefit from a community note. When enough requests are submitted, top writers can see alerts highlighting these posts and can choose to write a note in response (see the illustrations in Fig. 1). This division of roles—between note writers, note raters, and requestors—illustrates how Community Notes distributes responsibilities across different layers of the user base, thereby expanding participation beyond the core contributor group.
Research has shown that Community Notes system can successfully identify misleading content, reduce its spread, and even pressure authors to delete problematic posts (Wojcik et al., 2022; Martel and Rand, 2024; Allen et al., 2024b; Chuai et al., 2024a). Additionally, community-based annotations are often perceived as more trustworthy than professional fact-checks (Drolsbach et al., 2024). These encouraging findings suggest that Community Notes is promising in addressing misinformation on social media platforms and have sparked growing scholarly interest in understanding the dynamics of the system and identifying opportunities to refine and enhance its design (Augenstein et al., 2025). For instance, despite the relatively large volume of notes generated by contributors, only a small fraction—around 10%—achieve sufficient support from diverse users and are displayed to users (De et al., 2025; Chuai et al., 2024a). At the same time, the display process is often not fast enough to intervene before misleading posts have already begun to spread widely (Chuai et al., 2024a, b). To address these limitations, recent work has explored two complementary directions.
@@ 3.1. Data Collection
Subsequently, we used the full-archive search endpoint of the $\mathbb{X}$ Pro API to collect the requested posts. In total, the submitted requests target $2,427,451$ unique posts. However, according to the platform guidelines at the time of data collection, a post must receive at least 5 requests to be surfaced to top writers, indicating a threshold for which posts are eligible for community evaluation (Notes, 2025e). In our dataset, $154,090$ posts meet this criterion, representing only 6.3% of all requested posts. In contrast, the requests submitted to these posts account for more than half (52.6%) of the total requests. This suggests that, while users submitted a large volume of requests, attention was concentrated on a small set of posts. Based on the corresponding IDs of these eligible posts, we successfully retrieved $133,351$ posts, of which $98,685$ (74%) are in English and were authored by $22,915$ unique accounts. We restrict our analysis to the collected posts in English and focus on the request feature in the U.S. context. Next, we introduce how to extract the characteristics of posts and their authors for our subsequent analysis (see summary statistics in Table 1).
Table 1. Overview of requested posts. The three columns are for all requested posts, posts with community notes, and posts without community notes, respectively. The mean values for continuous variables (standard deviations in parentheses), percentages for dummy variables, or count numbers for count variables are reported.
@@
   Content types: Claim 0.687 (0.318) 0.687 (0.324) 0.687 (0.312)
   Content types: Opinion 0.489 (0.346) 0.475 (0.347) 0.506 (0.344)
   GPT misleadingness 0.398 (0.318) 0.405 (0.324) 0.389 (0.310)
@@
   Account types: Blue 72.5% 72.3% 72.6%
   Account types: Business 6.6% 6.4% 6.7%
   Account types: Government 7.2% 6.5% 7.9%